OpenStudy (rainbow_rocks03):

PLZ HELP AND EXPLAIN... Each side of a square is (7 + 3x) units. Which is the perimeter of the square?

OpenStudy (calculusxy):

a square has four sides and the perimeter is the total measurement of all of the side lengths. \[s + s + s + s = 4s \] \[4(7+ 3x) = ? \]

OpenStudy (calculusxy):

s = side length

OpenStudy (rainbow_rocks03):

So...(7+3x)=10 ?

OpenStudy (calculusxy):

no i meant for you to solve \(4(7+3x)\)

OpenStudy (rainbow_rocks03):

so like this...4*7=28+3x=31? IDK

OpenStudy (rainbow_rocks03):

@calculusxy ??

OpenStudy (rainbow_rocks03):

@calculusxy

OpenStudy (calculusxy):

\[4(7 - 3x) = 4(7) + 4(-3x) = 28 - 12x\]

OpenStudy (calculusxy):

you're distributing the four to the terms in the parenthesis

OpenStudy (calculusxy):

got it ?

OpenStudy (rainbow_rocks03):

ugh...no :(

OpenStudy (calculusxy):

where are you stuck?

OpenStudy (rainbow_rocks03):

I am stuck to where I get my answer.

OpenStudy (calculusxy):



OpenStudy (rainbow_rocks03):

So (7+3x) 4 times?So I multiply it 4 times?

OpenStudy (calculusxy):

yeah that's why it's \(4 \times (7 + 3x)\)

OpenStudy (rainbow_rocks03):

thats what i said!

OpenStudy (rainbow_rocks03):

look @calculusxy here are my answer choices...

OpenStudy (rainbow_rocks03):

A. 7 + 12x units, B. 28 + 3x units, C. 28 + 12x units, D. 49 + 9 units.